# Wireless 6G Connectivity for Massive Number of Devices and Critical Services

### v4

@@ Abstract
Compared to the generations up to 4G, whose main focus was on broadband and coverage aspects, 5G has expanded the scope of wireless cellular systems towards embracing two new types of connectivity: massive machine-type communication (mMTC) and ultra-reliable low-latency communications (URLLC). This paper will discuss the possible evolution of these two types of connectivity within the umbrella of 6G wireless systems. The paper consists of three parts. The first part deals with the connectivity for a massive number of devices. While mMTC research in 5G was predominantly focused on the problem of uncoordinated access in the uplink for a large number of devices, the traffic patterns in 6G may become more symmetric, leading to closed-loop massive connectivity. One of the drivers for this type of traffic patterns is distributed/decentralized learning and inference. The second part of the paper will discuss the evolution of wireless connectivity for critical services. While latency and reliability are tightly coupled in 5G, 6G will support a variety of safety critical control applications with different types of timing requirements, as evidenced by the emergence of metrics related to information freshness and information value. Additionally, ensuring ultra-high reliability for safety critical control applications requires modeling and estimation of the tail statistics of the wireless channel, queue length, and delay. The fulfillment of these stringent requirements calls for the development of novel AI-based techniques, incorporating optimization theory, explainable AI, generative AI and digital twins. The third part will analyze the coexistence of massive connectivity and critical services. Specifically, we will consider scenarios in which a massive number of devices need to support traffic patterns of mixed criticality. This will be followed by a discussion about the management of wireless resources shared by services with different criticality.
@@ I Introduction
One of the defining features of 5G was the native support for Internet of Things (IoT) through massive machine-type communication (mMTC) and ultra-reliable low-latency communications (URLLC). While mMTC aimed to provide delay-tolerant service for a very large number of low-cost devices, URLLC targeted human-centric and machine-centric real-time applications, such as virtual reality and self-driving cars. Along with enhanced mobile broadband (eMBB), these services constituted the “5G triangle,” which served as the guiding framework for the 5G research and standardization efforts [1]. At the conception of 5G, it had been thought that mMTC and URLLC, as well as their combinations with eMBB were sufficient to cover almost all IoT use cases. Nevertheless, with the evolution and deployment of 5G it started to emerge that the requirements for mMTC and URLLC would need to be redefined, as illustrated in Fig. 1. For instance, the stringent latency requirements of URLLC can be relaxed to a set of more general timing requirements that dynamically adjust to the current state of the system, such as Age of Information (AoI), Age of Loop (AoL), and Value of Information (VoI) [2]. Furthermore, new use cases, especially centered around machine learning (ML) and artificial intelligence (AI), have introduced new scenarios that were unthinkable when the 5G vision was initially sketched, such as federated learning and intelligent digital twins. Finally, there is a major change in the overall access infrastructure by embracing Non-Terrestrial Networks (NTN) [3, 4]. These networks introduce moving elements in the infrastructure, such as LEO satellites and significantly augment the spatial coverage, thereby the number of potentially accessing devices. Although NTNs will have an important role in massive connectivity, as also indicated in Fig. 1(b), in this article we will not cover the NTN aspects in detail, as it would require a more elaborate discussion on the specifics of the related network architecture, which is significantly different from the terrestrial one.
@@ II-C Sparse Recovery for Massive Random Access
As argued in the previous section, many modern random access methods rely on the use of non-orthogonal preambles to enlarge the set of available preambles, so as to reduce the probability of preamble collisions. However, the non-orthogonality of the transmitted preambles makes the activity detection problem nontrivial. The predominant strategy is to apply sparse recovery techniques that exploit the sparse user transmission patterns that result from the fact that only a small fraction of the users are active at any given time. In this section, we provide a brief review of the foundations behind this strategy in the context of massive random access. To simplify the presentation, we will focus on the case where each user is assigned a unique preamble, which precludes collisions. Clearly, supporting a massive number of users under this strategy is only possible with a very large number of preambles, which necessitates the preambles to be non-orthogonal. We consider a scenario with $N$ single-antenna devices and a single base station with $M$ antennas. Let $\mathbf{s}_{k}\in\mathbb{C}^{L\times 1}$ denote the preamble assigned to user $k$, $k=1,\ldots,N$, which is assumed to be known at the base station. Suppose a random number $K$ out of the $N$ users are active, and let $a_{k}=1$ if user $k$ is active, otherwise $a_{k}=0$. The channel between device $k$ and the base station is assumed to be given by the random vector $g_{k}\mathbf{h}_{k}\in\mathbb{C}^{M\times 1}$, composed by large scale fading coefficient $g_{k}\in\mathbb{R}$, and unknown Rayleigh fading coefficients $\mathbf{h}_{k}\in\mathbb{C}^{M}$ drawn independently from $\mathcal{CN}(0,1)$. The signal $\mathbf{Y}\in\mathbb{C}^{L\times M}$ received by the base station can then be written as
@@ II-D2 Transmission of General Feedback Messages
We consider next the problem of transmitting $K$ general messages $W_{1^{\prime}},W_{2^{\prime}},\ldots,W_{K^{\prime}}\in\mathcal{M}$ to $K$ out of the $N$ users, which was recently studied in [35, 36]. Compared to the previous case, instead of assuming a false positive rate, we will assume that the devices know whether their own message is informative or empty, but have no information about the other messages. This assumption is often reasonable, e.g., if the messages are transmitted as responses to an uplink transmission. Alternatively, the message acknowledgment scheme presented in the previous section can be used to indicate the intended recipients. The central problem is then to encode the $K$ messages in such a way that the recipients know how to decode their own message without knowing the identities of the $K-1$ other recipients of informative messages. Note that, if the users know perfectly the identifiers of all $K$ recipients, this can be done simply by listing the messages in the order of the identifiers. However, when each user only knows whether there is a packet for them or not, and not who the other recipients are, the problem becomes more involved. Assuming that the messages are independent and identically distributed across users, the naive encoding scheme of concatenating each of the $K$ messages with the identity of its recipient requires approximately $kH(W_{n})+K\log_{2}N$ bits, where $H(W_{n})$ denotes the entropy of a message $W_{n}$. However as shown in [35], theoretically speaking, the overhead of $K\log_{2}N$ bits can in fact be reduced to only $O(\log_{2}K)$ bits for any independent and identically distributed (i.i.d.) messages, and further to $O(1)$ when the messages are equally likely across the alphabet, thus removing the dependency on $N$.
@@ II-E Integration of Massive Connectivity in 6G Standardization
In the context of downlink, the techniques discussed in the previous sections are more straightforward to realize, as they do not require modifications to the physical layer. For the same reason, they do not rely on specific assumptions that need to be validated experimentally. However, while it may be unlikely that downlink designs tailored for specific applications will be implemented in early 6G, the theoretical insights from Section II-D provide useful guidelines in the design of the control information for massive access in 6G, such as the design of acknowledgments and scheduling for grant-based access.
@@ III-C1 Network Design for Control Systems
  • Stochastic Maximum Allowed Transfer Interval (MATI) and Maximum Allowed Delay (MAD): The criteria governing the performance and stability of control systems are expressed through the definition of two key parameters: Maximum Allowed Transfer Interval (MATI) and Maximum Allowable Delay (MAD) [51, 52]. While wireline networks can often meet these rigorous real-time criteria, the same expectations are infeasible in wireless networks due to inherent packet error probability. Consequently, wireless control systems adopt a stochastic MATI constraint, ensuring that the time interval between successive state vector reports remains above MATI value with a predefined probability [53, 54].
@@ III-D AI Based Techniques for Critical Services
  • Optimization theory based AI: Optimization theory based AI approach aims to leverage optimization theory knowledge to improve the performance of AI algorithms. This reduces the reliance on extensive training data, leading to minimized delays and improved solution reliability and robustness [89, 90, 56, 91]. In [89], a deep neural network is first pretrained using synthetic data generated from an optimization-theory based algorithm, followed by fine-tuning in a real environment with actual measured data. In DNN, system parameters are mapped to optimal resource allocation. Pretraining DNN involves generating a diverse set of system configurations, representing all possible inputs, and computing the optimal resource allocation using the optimization-theory based algorithm. This pretraining substantially decreases the amount of real-time data required for the implementation of AI approaches.
@@ IV-A From Network Slicing to Intent-Based Management
Orchestrating and managing the different network slices can be a complex task [99, 100]. It is expected that 6G will leverage the use ML techniques to deploy intent-based network-management solutions [101]. The underlying idea is that network management will be performed via an abstraction layer, which will interpret and automatically translate the intents into network configurations and, while operating the network, continuously learn and fine-tune the configurations based on the ongoing data communication. Given the recent advances in natural language processing, future systems may even allow the intents to be given in human language (e.g., “configure the network to enable $50$ IoT devices to share their local machine-learning models via federated learning, with a high differential-privacy requirement”). Intent-based network management is a powerful tool that could enable an operator to offer service-layer agreements beyond the best-effort type of services provided today. For example, it could allow an operator to guarantee a certain data rate and latency characteristic throughout a city as the intent-based management automatically and continuously monitors and adjusts the network configuration to meet this requirement, possibly by prioritizing this traffic over other traffic types such as massive IoT connections, which may have significantly relaxed latency requirements in comparison.
@@ IV-B Slicing of RAN Communication Resources: A Case Study
Let us first focus on the coexistence between broadband traffic and critical services as originally defined in 5G. From a communication theory perspective, we can model the broadband-traffic transmission problem as a problem of communicating information over multiple frequency channels (see Fig. 5(a)) over a fading channel that can be assumed to be known to both transmitter and receiver. This assumption is a plausible approximation, because of the large payloads that are typical of broadband traffic, which make the cost of exchanging channel-state information negligible. One reasonable objective when transmitting broadband traffic is to maximize the transmission rate under a long-term power constraint as well as a constraint on the outage probability.

### v5

@@ Abstract
Compared to the generations up to 4G, whose main focus was on broadband and coverage aspects, 5G has expanded the scope of wireless cellular systems towards embracing two new types of connectivity: massive machine-type communication (mMTC) and ultra-reliable low-latency communications (URLLC). This paper discusses the possible evolution of these two types of connectivity within the umbrella of 6G wireless systems. The paper consists of three parts. The first part deals with the connectivity for a massive number of devices. While mMTC research in 5G predominantly focuses on the problem of uncoordinated access in the uplink for a large number of devices, the traffic patterns in 6G may become more symmetric, leading to closed-loop massive connectivity. One of the drivers for this type of traffic patterns is distributed/decentralized learning and inference. The second part of the paper discusses the evolution of wireless connectivity for critical services. While latency and reliability are tightly coupled in 5G, 6G will support a variety of safety critical control applications with different types of timing requirements, as evidenced by the emergence of metrics related to information freshness and information value. Additionally, ensuring ultra-high reliability for safety critical control applications requires modeling and estimation of the tail statistics of the wireless channel, queue length, and delay. The fulfillment of these stringent requirements calls for the development of novel AI-based techniques, incorporating optimization theory, explainable AI, generative AI and digital twins. The third part analyzes the coexistence of massive connectivity and critical services. Specifically, we consider scenarios in which a massive number of devices need to support traffic patterns of mixed criticality. This is followed by a discussion about the management of wireless resources shared by services with different criticality.
@@ I Introduction
One of the defining features of 5G was the native support for Internet of Things (IoT) through massive machine-type communication (mMTC) and ultra-reliable low-latency communications (URLLC). While mMTC aimed to provide delay-tolerant service for a very large number of low-cost devices, URLLC targeted human-centric and machine-centric real-time applications, such as virtual reality and self-driving cars. Along with enhanced mobile broadband (eMBB), these services constituted the “5G triangle,” which served as the guiding framework for the 5G research and standardization efforts [1]. At the conception of 5G, it had been thought that mMTC and URLLC, as well as their combinations with eMBB were sufficient to cover almost all IoT use cases. Nevertheless, with the evolution and deployment of 5G it started to emerge that the requirements for mMTC and URLLC would need to be redefined, as illustrated in Fig. 1. For instance, the stringent latency requirements of URLLC can be relaxed to a set of more general timing requirements that dynamically adjust to the current state of the system, such as Age of Information (AoI), Age of Loop (AoL), and Value of Information (VoI) [2]. Furthermore, new use cases, especially centered around machine learning (ML) and artificial intelligence (AI), have introduced new scenarios that were unthinkable when the 5G vision was initially sketched, such as federated learning and intelligent digital twins. Finally, there is a major change in the overall access infrastructure by embracing nonterrestrial networks (NTNs) [3, 4]. These networks introduce moving elements in the infrastructure, such as low Earth orbit (LEO) satellites and significantly augment the spatial coverage, thereby the number of potentially accessing devices. Although NTNs will have an important role in massive connectivity, as also indicated in Fig. 1(b), in this article we will not cover the NTN aspects in detail, as it would require a more elaborate discussion on the specifics of the related network architecture, which is significantly different from the terrestrial one.
@@ II-C Sparse Recovery for Massive Random Access
As argued in Section II-B, many modern random access methods rely on the use of non-orthogonal preambles to enlarge the set of available preambles, so as to reduce the probability of preamble collisions. However, the non-orthogonality of the transmitted preambles makes the activity detection problem nontrivial. The predominant strategy is to apply sparse recovery techniques that exploit the sparse user transmission patterns that result from the fact that only a small fraction of the users are active at any given time. In this section, we provide a brief review of the foundations behind this strategy in the context of massive random access. To simplify the presentation, we will focus on the case where each user is assigned a unique preamble, which precludes collisions. Clearly, supporting a massive number of users under this strategy is only possible with a very large number of preambles, which necessitates the preambles to be non-orthogonal. We consider a scenario with $N$ single-antenna devices and a single base station with $M$ antennas. Let $\mathbf{s}_{k}\in\mathbb{C}^{L\times 1}$ denote the preamble assigned to user $k$, $k=1,\ldots,N$, which is assumed to be known at the base station. Suppose a random number $K$ out of the $N$ users are active, and let $a_{k}=1$ if user $k$ is active, otherwise $a_{k}=0$. The channel between device $k$ and the base station is assumed to be given by the random vector $g_{k}\mathbf{h}_{k}\in\mathbb{C}^{M\times 1}$, composed by large scale fading coefficient $g_{k}\in\mathbb{R}$, and unknown Rayleigh fading coefficients $\mathbf{h}_{k}\in\mathbb{C}^{M}$ drawn independently from $\mathcal{CN}(0,1)$. The signal $\mathbf{Y}\in\mathbb{C}^{L\times M}$ received by the base station can then be written as
@@ II-D2 Transmission of General Feedback Messages
We consider next the problem of transmitting $K$ general messages $W_{1^{\prime}},W_{2^{\prime}},\ldots,W_{K^{\prime}}\in\mathcal{M}$ to $K$ out of the $N$ users, which was recently studied in [35, 36]. Compared to the previous case, instead of assuming a false positive rate, we will assume that the devices know whether their own message is informative or empty, but have no information about the other messages. This assumption is often reasonable, e.g., if the messages are transmitted as responses to an uplink transmission. Alternatively, the message acknowledgment scheme presented in Section II-D1 can be used to indicate the intended recipients. The central problem is then to encode the $K$ messages in such a way that the recipients know how to decode their own message without knowing the identities of the $K-1$ other recipients of informative messages. Note that, if the users know perfectly the identifiers of all $K$ recipients, this can be done simply by listing the messages in the order of the identifiers. However, when each user only knows whether there is a packet for them or not, and not who the other recipients are, the problem becomes more involved. Assuming that the messages are independent and identically distributed across users, the naive encoding scheme of concatenating each of the $K$ messages with the identity of its recipient requires approximately $kH(W_{n})+K\log_{2}N$ bits, where $H(W_{n})$ denotes the entropy of a message $W_{n}$. However as shown in [35], theoretically speaking, the overhead of $K\log_{2}N$ bits can in fact be reduced to only $O(\log_{2}K)$ bits for any independent and identically distributed (i.i.d.) messages, and further to $O(1)$ when the messages are equally likely across the alphabet, thus removing the dependency on $N$.
@@ II-E Integration of Massive Connectivity in 6G Standardization
In the context of downlink, the techniques discussed in Section II-D are more straightforward to realize, as they do not require modifications to the physical layer. For the same reason, they do not rely on specific assumptions that need to be validated experimentally. However, while it may be unlikely that downlink designs tailored for specific applications will be implemented in early 6G, the theoretical insights from Section II-D provide useful guidelines in the design of the control information for massive access in 6G, such as the design of acknowledgments and scheduling for grant-based access.
@@ III-C1 Network Design for Control Systems
  • Stochastic maximum allowable transfer interval (MATI) and maximum allowable delay (MAD): The criteria governing the performance and stability of control systems are expressed through the definition of two key parameters: MATI and MAD [51, 52]. While wireline networks can often meet these rigorous real-time criteria, the same expectations are infeasible in wireless networks due to inherent packet error probability. Consequently, wireless control systems adopt a stochastic MATI constraint, ensuring that the time interval between successive state vector reports remains above MATI value with a predefined probability [53, 54].
@@ III-D AI Based Techniques for Critical Services
  • Optimization theory based AI: Optimization theory based AI approach aims to leverage optimization theory knowledge to improve the performance of AI algorithms. This reduces the reliance on extensive training data, leading to minimized delays and improved solution reliability and robustness [89, 90, 56, 91]. In [89], a deep neural network (DNN) is first pretrained using synthetic data generated from an optimization-theory based algorithm, followed by fine-tuning in a real environment with actual measured data. In DNN, system parameters are mapped to optimal resource allocation. Pretraining DNN involves generating a diverse set of system configurations, representing all possible inputs, and computing the optimal resource allocation using the optimization-theory based algorithm. This pretraining substantially decreases the amount of real-time data required for the implementation of AI approaches.
@@ IV-A From Network Slicing to Intent-Based Management
Orchestrating and managing the different network slices can be a complex task [99, 100]. It is expected that 6G will leverage the use ML techniques to deploy intent-based network-management solutions [101]. The underlying idea is that network management will be performed via an abstraction layer, which will interpret and automatically translate the intents into network configurations and, while operating the network, continuously learn and fine-tune the configurations based on the ongoing data communication. Given the recent advances in natural language processing, future systems may even allow the intents to be given in human language (e.g., “configure the network to enable $50$ IoT devices to share their local machine-learning models via federated learning, with a high differential-privacy requirement”). Intent-based network management is a powerful tool that could enable an operator to offer service-level agreements beyond the best-effort type of services provided today. For example, it could allow an operator to guarantee a certain data rate and latency characteristic throughout a city as the intent-based management automatically and continuously monitors and adjusts the network configuration to meet this requirement, possibly by prioritizing this traffic over other traffic types such as massive IoT connections, which may have significantly relaxed latency requirements in comparison.
@@ IV-B Slicing of RAN Communication Resources: A Case Study
Let us first focus on the coexistence between broadband traffic and critical services as originally defined in 5G. From a communication theory perspective, we can model the broadband-traffic transmission problem as a problem of communicating information over multiple frequency channels (see Fig. 5(a)) over a fading channel that can be assumed to be known to both transmitter and receiver. This assumption is a plausible approximation, because of the large payloads that are typically assumed for broadband traffic, which make the cost of exchanging channel-state information negligible. One reasonable objective when transmitting broadband traffic is to maximize the transmission rate under a long-term power constraint as well as a constraint on the outage probability.